OpenStudy (anonymous):

Simplify radical 64.

OpenStudy (anonymous):

8

OpenStudy (anonymous):

since \[8^2=64\]

OpenStudy (anonymous):

Simplifying and solving are two different things, isnt it?

OpenStudy (anonymous):

\[64=8^2\]So\[\sqrt{64} = \sqrt{8^2} = 8\] Solving generally requires an equation with one or more variables.

OpenStudy (anonymous):

simplifying just means 'rewrite' in the form I want.

OpenStudy (anonymous):

well unfortunately many math teachers use "simply" to mean all kinds of different things, confusing the hell out students

OpenStudy (anonymous):

Where the form they're looking for isn't always clear.

OpenStudy (anonymous):

there is no such math term as "simplify" it is pure laziness or confusion on their part

OpenStudy (anonymous):

sometimes it just means compute

OpenStudy (anonymous):

sometimes it means multiply sometimes it means combine like terms sometimes it means factor and cancel you never know

OpenStudy (anonymous):

nothing on earth is more confusing to use one term or one symbol to mean many different things. sheer utter laziness is what it is, usually from people who should not be teaching math or writing books

OpenStudy (anonymous):

Tell me about it... hahah . I thought you had to do something like radical 2*2*2*2*2*2 and then pair the twos and then take them out, so you would get a normal 6 instead of a radical 8.. But i dont know what they mean by simplify. :l

OpenStudy (anonymous):

there is such a thing as "simplest radical form" and it has specific rules in this case you just have the square root of 64 which is pretty obviously 8. only reason you thought twice was because some moron wrote "simplify"

OpenStudy (anonymous):

You can certainly do that if you like: \[64 = 2\times 2 \times 2 \times 2\times 2 \times 2\]So \[\sqrt{64} = \sqrt{2\times 2 \times 2 \times 2\times 2 \times 2} = 2\times 2\times 2 = 8 \]

OpenStudy (anonymous):

You take out a pair of 2's and it becomes 1 two after you take the square root.

OpenStudy (anonymous):

But that seems too easy... I mean, its geomerty homework! Shouldnt it be all confusing and such? and ohh! i got it! i did my math wrong. It all does the same thing...

OpenStudy (anonymous):

sure if you have infinite patience. i guess you could do that with \[\sqrt{169}\] or \[\sqrt{1,000,000}\] if you like

OpenStudy (anonymous):

Mkay, so how do you multiply radicals?

OpenStudy (anonymous):

Lol, no it should be easy. =)

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{b} = \sqrt{a\times b}\]

OpenStudy (anonymous):

if the index is the same multiply the radicands. what polpak said

OpenStudy (anonymous):

Oh.. okay. uh, what about radical 40? Cause that cant be squarre rooted like 64...

OpenStudy (anonymous):

That's ok, it can still be factored.

OpenStudy (anonymous):

into what?

OpenStudy (anonymous):

\[40=4\times 10\] \[\sqrt{40}=\sqrt{4\times 10}=\sqrt{4}\times \sqrt{10}=2\sqrt{10}\]

OpenStudy (anonymous):

the it is in all gory details

OpenStudy (anonymous):

Mkay, thats what i thought....

OpenStudy (anonymous):

not too bad. factor the number and make sure to take out all perfect squares if it is a square root

OpenStudy (karatechopper):

8 times 8 is 64, so 8

OpenStudy (anonymous):

what about dividing radicals?

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

\[\sqrt{\frac{64}{49}}=\frac{8}{7}\] for example

OpenStudy (anonymous):

\[40 = 4\times 10\] \[\implies \sqrt{40} = \sqrt{4\times 10} = \sqrt{4}\times \sqrt{10}\]

OpenStudy (anonymous):

bah, little late on that reply ;p

OpenStudy (anonymous):

in other words \[\sqrt{\frac{a}{b}}=\frac{\sqrt{a}}{\sqrt{b}}\]

OpenStudy (anonymous):

@polpak not as late as karatechopper!

OpenStudy (anonymous):

So you can do each problem sepratley, and then then just put the line thing in between them and you have your answer?

OpenStudy (karatechopper):

hey hey hey i just logged on to openstudy! whats my fault in it?

OpenStudy (anonymous):

well you have a choice. you can divide and take the square root second or take the square root separately and then divide. the equal sign means you get the same thing in either case

OpenStudy (anonymous):

@karatechopper no offense meant. sorry

OpenStudy (anonymous):

you have an example?

OpenStudy (anonymous):

radical 120 over radical 10

OpenStudy (anonymous):

so you can divide 120 by 10 to get 12 first

OpenStudy (anonymous):

then you would have \[\sqrt{12}\]

OpenStudy (anonymous):

and since \[12=4\times 3\] you get \[\sqrt{12}=\sqrt{4\times 3}=2\sqrt{3}\]

OpenStudy (karatechopper):

apology exepted i will become ur fan can u become mine?

OpenStudy (anonymous):

Mkay, that makes sense. What about if you have a problem like 3 radical 5 * 2 radical 10? what do you do to the numbers that arent in the radical?

OpenStudy (anonymous):

make them the same

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

\[\sqrt[3]{5}\sqrt{10}\] is that it?

OpenStudy (anonymous):

Typically you'd do that with fractional exponents.

OpenStudy (anonymous):

unfortunately you have to make them in to 6ht roots since that is the least common multiple of 2 and 3

OpenStudy (anonymous):

Converting your radical into a fractional exponent.

OpenStudy (anonymous):

for example \[\sqrt{10}=\sqrt[6]{10^3}\]

OpenStudy (anonymous):

and \[\sqrt[3]{5}=\sqrt[6]{5^2}\]

OpenStudy (anonymous):

so now that the indexes are the same you can combine them

OpenStudy (anonymous):

... Im confused. why do you have to change them? Whats an index?

OpenStudy (anonymous):

and get \[\sqrt[6]{5^2\times 10^3}\]

OpenStudy (anonymous):

ok lets to slow

OpenStudy (anonymous):

Have you learned about fractional exponents and their relationship to radicals?

OpenStudy (anonymous):

\[\sqrt{5} = 5^{\frac{1}{2}}\]

OpenStudy (anonymous):

the index is the number outside the radical. so for example in \[\sqrt[5]{x^8}\] the index is 5

OpenStudy (anonymous):

this means the fifth root

OpenStudy (anonymous):

in \[\sqrt[3]{5}\] the index is 3. this means the cube root

OpenStudy (anonymous):

if you do not see an index, for example in \[\sqrt{5}\] the index is assumed to be 2

OpenStudy (anonymous):

Uh... Probbably. I just dont remember anything. Math was my first class of the day, and you cant really learn much at 7 in the monring. And ... blah! this is still confusing. This is what my equation looks like.

OpenStudy (anonymous):

^^number 5.

OpenStudy (anonymous):

oh my dear. sorry for confusing you

OpenStudy (anonymous):

this is not an index, this is just a number.

OpenStudy (anonymous):

Thats what i thougght!(:

OpenStudy (anonymous):

answer to number 5 is \[6\sqrt{50}\]

OpenStudy (anonymous):

easy easy easy

OpenStudy (anonymous):

Is that simplified?

OpenStudy (anonymous):

right? \[2\times 3=6\] \[5\times 10=50\]

OpenStudy (anonymous):

no now we simplify. \[50=25\times 2\] so \[\sqrt{50}=5\sqrt{2}\]

OpenStudy (anonymous):

because \[\sqrt{50}=\sqrt{25\times 2}=5\sqrt{2}\]

OpenStudy (anonymous):

so you get \[6\times 5\times \sqrt{2}=30\sqrt{2}\]

OpenStudy (anonymous):

Oh... Wait what about the 6? The final answer would be 30 radical 2 right?

OpenStudy (anonymous):

\[6\sqrt{50} \]\[= 6 \times \sqrt{50}\]\[=6 \times \sqrt{5\times 5 \times 2}\]\[ = 6\times 5\times \sqrt{2}\]\[=30\sqrt{2}\]

OpenStudy (anonymous):

Alright, it makes sense. Much easier than all that index jazz. (:

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i misread what you wrote. sorry

OpenStudy (anonymous):

Yes, indexes will come later ;)

OpenStudy (anonymous):

Not that they're harder, but you have to be able to do this well before you can do that well ;)

OpenStudy (anonymous):

Hah, its alright . How do you subtract radicals? like... 3 radical 6 - 8 radical 6.?

OpenStudy (anonymous):

You cannot. You have to make them products of similar factors, then factor out what's in common.

OpenStudy (anonymous):

So \[3\sqrt{6} - 8\sqrt{6} = (3-8)\sqrt{6}\]

OpenStudy (anonymous):

Actually this one you can think of like oranges. If you have 3 oranges and you subtract 8 oranges you will have how many oranges? If you have 3 sqrt(6)s and you subtract 8 sqrt(6)s how many sqrt(6)s will you have left?

OpenStudy (anonymous):

... uh. none?

OpenStudy (anonymous):

You'll have -5

OpenStudy (anonymous):

Ok, the oranges things breaks down a bit when talking about negative numbers, but you get the idea ;p

OpenStudy (anonymous):

So... the answer is -5?

OpenStudy (anonymous):

it's \[-5\sqrt{6}\]

OpenStudy (anonymous):

If it helps you to think about it: \[let\ a = \sqrt{6}\] \[3\sqrt{6} - 8\sqrt{6} = 3a - 8a = -5a = -5 \sqrt{6}\]

OpenStudy (anonymous):

Oh allright, thank you. (: Onne more question, how do you divide radical 50 over radical 49?

OpenStudy (anonymous):

\[\frac{\sqrt{50}}{\sqrt{49}} = \frac{\sqrt{50}}{7} = \frac{\sqrt{5\times 5\times 2}}{7} = \frac{5\sqrt{2}}{7}\]